OpenStudy (anonymous):

Verify the applicability of the Intermediate Value Theorem (IVT) in the interval [0,5] and find the value c guaranteed by the theorem. f(x)= x^2 + x - 1, f(c) = 11

OpenStudy (anonymous):

i can tell you right off the bat that c is smack dab in the middle because you have a quadratic. we can find it if you like

OpenStudy (anonymous):

please explain!

OpenStudy (anonymous):

ok what is \[f(5)\]?

OpenStudy (anonymous):

29

OpenStudy (anonymous):

yes and of course \[f(0)=-1\]

OpenStudy (anonymous):

so this graph contains the points \[(0,-1),(5,29)\] and the slope between those two points is \[\frac{29+1}{5}=\frac{30}{5}=6\]

OpenStudy (zarkon):

?

OpenStudy (anonymous):

f(0) = -1 f(5) = 29 and 11 is betwen those so IVT applies

OpenStudy (anonymous):

to find c you x^2 +x -1 = 11

OpenStudy (anonymous):

lorda mercy it says "intermediate value theorem" and i read "mean value theorem" i should learn to read

OpenStudy (zarkon):

;)

OpenStudy (anonymous):

thank you all!

OpenStudy (anonymous):

i did this exact problem in my class

OpenStudy (anonymous):

in that case solve \[x^2+x-1=11\]which must have a solution because \[-1<11<29\]

OpenStudy (anonymous):

is that it?

OpenStudy (zarkon):

we should also mention that the function is a polynomial and is therefore continuous on the interval we are interested in